The string of pendulum of length l is displaced through 90 degree from the vertical and released .Then the minimum strength of the string in order to withstand the tension as the pendulum passes through the mean position is

Grade:12

4 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

when pendulam is reliesed them decrese in potential energy will increase the kinetic energy

energy balance

 mg(l/2)  = mv2/2

 v2= gl ..........1

now at lower most position centrifugal force as well as weight both acting in downward direction which is balance by tension in string

so T = mg + mv2/(l/2)

       = mg + 2mg

         =3mg

        

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

 

sushama shinde
6 years ago
Enter your answer here...Velocity at lowest point v= root2gl......(1) At the lowest point the tension in the string T=mg+mv2/l........ (2) put eq.1 in 2 T=mg+m2gl/l = 3mg
rishi
6 years ago
2 stone projected with the same speed but the different but making different angle with horizontal their original drainage equal the angle of projection of 1 is 5 by 3 and the maximum height reached by it is 0.2 the maximum height reached by the other is meta
rishi
6 years ago
a simple pendulum of mass 2200 gram and length is length 1m is turn 8 to 60 degree and then release event pendulum makes an angle of 30 degree with the mean position its kinetic and potential velocity is